OpenStudy (anonymous):

\[-\frac{ 5 }{ 21 }-\frac{ 4\sqrt{23} }{ 21 }=??????\]

OpenStudy (ash2326):

it's just \[\frac {-5-4\sqrt {23}}{21}\] or \[-\frac {5+4\sqrt {23}}{21}\] if you wanna simplify it more you'd have to find value of \(\sqrt{23}\) and then find the value

OpenStudy (ash2326):

*of the expression

OpenStudy (anonymous):

oh ok sweet